OpenStudy (anonymous):

How many times does the graph of y = 4x^2-12x+9 intersect the x-axis?

OpenStudy (anonymous):

when the graph intersects the xaxis then y=0 4x^2-12x+9=0 (2x-3)^2=0 2x-3=0 x=3/2 so the point is (3/2,0)